# Non-Rigid Designators in Modal and Temporal Free Description Logics

### v1

@@ Appendix D Proofs for Section 5
    $W^{\prime}=(W\setminus W_{\vec{w}})\cup\{\vec{w}^{\prime}i\vec{r}\mid\vec{v}^{\prime}j\vec{r}\in W_{\vec{v}}\}.$
  • Set $\boldsymbol{q}^{\prime}(\vec{u})=\boldsymbol{q}(\vec{u})$ for all $\vec{u}\in W\setminus W_{\vec{w}}$ and $\boldsymbol{q}^{\prime}(\vec{w}^{\prime}i\vec{r})=\boldsymbol{q}(\vec{v}^{\prime}j\vec{r})$ for all $\vec{v}^{\prime}j\vec{r}\in W_{\vec{v}}$.
  • Take for any $\rho\in\mathfrak{R}$ with $\vec{w}^{\prime}\in\text{dom}(\rho)$ a $f(\rho)\in\mathfrak{R}$ with $\vec{v}\in\text{dom}(f(\rho))$ such that $f(\rho)(\vec{v})=\rho(\vec{w})$. We can ensure that all $f(\rho)$ are distinct since $x_{i}\leq y_{i}$ for all $i\leq k$. Then let $\rho^{\prime}\in\mathfrak{R}^{\prime}$ with $\rho^{\prime}$ defined by setting $\rho^{\prime}(\vec{u})=\rho(\vec{u})$ for $\vec{u}\in W\setminus W_{\vec{w}}$ and $\rho^{\prime}(\vec{w}^{\prime}i\vec{r})=f(\rho)(\vec{v}^{\prime}j\vec{r})$ for all $\vec{v}^{\prime}j\vec{r}\in W_{\vec{v}}$.
    Also add all $\rho\in\mathfrak{R}$ with $\text{dom}(\rho)\cap W_{\vec{w}}=\emptyset$ to $\mathfrak{R}^{\prime}$. Finally, for all $\rho$ with $\text{dom}(\rho)\subseteq W_{\vec{w}}$ which were not selected as an $f(\rho)$, add the $\rho^{\prime}$ to $\mathfrak{R}^{\prime}$ defined by setting $\rho^{\prime}(\vec{w}^{\prime}i\vec{r})=\rho(\vec{v}^{\prime}j\vec{r})$ for all $\vec{v}^{\prime}j\vec{r}\in W_{\vec{v}}$.
@@
Drop interval 2: Consider the root $w_{0}$ and $v\in W$ with
  • $\boldsymbol{q}(w_{0})=(x_{1},\ldots,x_{k})$;
  • $\boldsymbol{q}(v)=(y_{1},\ldots,y_{k})$;
  • $x_{i}\leq y_{i}$ for all $i\leq k$.
Then construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by restricting $\mathfrak{Q}$ to $W_{v}$ in the obvious way.
Restrict outdegree. Consider $\vec{w}\in W$ and assume $\boldsymbol{q}(\vec{w})=(x_{1},\ldots,x_{k})$. To bound the number of successors of $\vec{w}$ to $|\boldsymbol{q}(\vec{w})|\times|C_{0}|$ pick for every $j\leq k$ with $x_{j}>0$ and every $\Diamond_{i}C\in{\boldsymbol{t}}_{j}$ for every run $\rho\in\mathfrak{R}$ with $\rho(\vec{w})=j$ a successor $\vec{w}iw\in W$ with $C\in{\boldsymbol{t}}_{\rho(\vec{w}iw)}$ or $\Diamond_{i}C\in{\boldsymbol{t}}_{\rho(\vec{w}iw)}$. Let $V$ be the set of successors of $W$ picked in this way. Then construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by restricting $\mathfrak{Q}$ to
$W^{\prime}=(W\setminus W_{\vec{w}})\cup\{\vec{w}\}\cup\bigcup_{\vec{v}\in V}W_{\vec{v}}.$
It is easy to see that $\mathfrak{Q}^{\prime}$ is a quasimodel for $C_{0}$ such that the outdegree of $\vec{w}$ is bounded by $|\boldsymbol{q}(\vec{w})|\times|C_{0}|$.
@@
As argued above, it follows from Lemma 27 that we have a recursive bound on the size of a quasimodel for $C_{0}$. By Lemma 26, we then also have a recursive bound on the size of a finite tree-shaped model satisfying $C_{0}$. This shows the decidability of concept-satisfiability for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$ in expanding domain models.
We now discuss concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ and $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ under expanding domains. We focus on $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$, $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ can be treated similarly. Note that a straightforward selective filtration argument applied to appropriate quasimodels shows that in $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ over expanding domain models every satisfiable concept is satisfied in a finite model. This observation implies that a $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}}$-concept $C$ with a single modal operator is satisfiable in an expanding domain model for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ iff it is satisfiable in a model for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ with expanding domains, where the modal operator in $C$ is interpreted by the transitive closure of $R_{1}\cup\cdots\cup R_{n}$. Hence the decidability of concept-satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ over expanding domain models is a consequence of the decidability of concept satisfiability for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ over expanding domain models.
The decidability of concept-satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ over expanding domain models can also be derived from decidability results for expanding domain product modal logics. We define the relevant modal logics using DL syntax. Consider the modal DL with a single role, $o$, interpreted as a Noetherian strict linear order. In DL syntax, concepts in $\mathcal{ML}^{n}_{lin}$ are of the form
@@
The proof is similar to the proof of Theorem 10 (1). Assume an $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concept $C$ is given. Assume for simplicity that $C$ contains a single modal operator, denoted by $\Box$. Let $m$ be the modal depth of $C$. Let $\mathsf{con}(C)$ be the closure under single negation of the set of concepts occurring in $C$. A type for $C$ is a subset ${\boldsymbol{t}}$ of $\mathsf{con}(C)$ such that $\neg C\in{\boldsymbol{t}}$ iff $C\not\in{\boldsymbol{t}}$, for all $\neg C\in\mathsf{con}(\mathcal{O})$. A quasistate for $D$ is a non-empty set $\boldsymbol{T}$ of types for $C$. The description of $\boldsymbol{T}$ is the $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concept
@@
Now decidability of concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ over expanding domain models follows from decidability of satisfiability of $\mathcal{ML}^{1}_{lin}$-concepts in models with expanding domains and transitive Noetherian frames (?).

### v2

@@ D.2 Proof of Theorem 12
    $W^{\prime}=(W\setminus W_{\boldsymbol{w}})\cup\{\boldsymbol{w}^{\prime}i\boldsymbol{r}\mid\boldsymbol{v}^{\prime}j\boldsymbol{r}\in W_{\boldsymbol{v}}\}.$
  • Set $\boldsymbol{q}^{\prime}(\boldsymbol{u})=\boldsymbol{q}(\boldsymbol{u})$ for all $\boldsymbol{u}\in W\setminus W_{\boldsymbol{w}}$ and $\boldsymbol{q}^{\prime}(\boldsymbol{w}^{\prime}i\boldsymbol{r})=\boldsymbol{q}(\boldsymbol{v}^{\prime}j\boldsymbol{r})$ for all $\boldsymbol{v}^{\prime}j\boldsymbol{r}\in W_{\boldsymbol{v}}$.
  • Take, for any $\rho\in\mathfrak{R}$ with $\boldsymbol{w}^{\prime}\in\text{dom}(\rho)$, an $f(\rho)\in\mathfrak{R}$ with $\boldsymbol{v}\in\text{dom}(f(\rho))$ such that $f(\rho)(\boldsymbol{v})=\rho(\boldsymbol{w})$. We can ensure that all $f(\rho)$ are distinct since $\boldsymbol{q}(\boldsymbol{w})\leq\boldsymbol{q}(\boldsymbol{v})$. Then let $\rho^{\prime}\in\mathfrak{R}^{\prime}$ with $\rho^{\prime}$ be defined by setting $\rho^{\prime}(\boldsymbol{u})=\rho(\boldsymbol{u})$ for $\boldsymbol{u}\in W\setminus W_{\boldsymbol{w}}$ and $\rho^{\prime}(\boldsymbol{w}^{\prime}i\boldsymbol{r})=f(\rho)(\boldsymbol{v}^{\prime}j\boldsymbol{r})$ for all $\boldsymbol{v}^{\prime}j\boldsymbol{r}\in W_{\boldsymbol{v}}$.
    Also add all $\rho\in\mathfrak{R}$ with $\text{dom}(\rho)\cap W_{\boldsymbol{w}}=\emptyset$ to $\mathfrak{R}^{\prime}$. Finally, for all $\rho$ with $\text{dom}(\rho)\subseteq W_{\boldsymbol{w}}$ which were not selected as an $f(\rho)$, add the $\rho^{\prime}$ to $\mathfrak{R}^{\prime}$ defined by setting $\rho^{\prime}(\boldsymbol{w}^{\prime}i\boldsymbol{r})=\rho(\boldsymbol{v}^{\prime}j\boldsymbol{r})$ for all $\boldsymbol{v}^{\prime}j\boldsymbol{r}\in W_{\boldsymbol{v}}$.
@@
Drop interval 2: Consider the root $w_{0}$ and $\boldsymbol{v}\in W$ with $\boldsymbol{q}(w_{0})\leq\boldsymbol{q}(\boldsymbol{v})$. Then construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by restricting $\mathfrak{Q}$ to $W_{\boldsymbol{v}}$ in the obvious way.
Restrict outdegree. Consider $\boldsymbol{w}\in W$ and assume $\boldsymbol{q}(\boldsymbol{w})=(x_{1},\ldots,x_{k})$. To bound the number of successors of $\boldsymbol{w}$ to $|\boldsymbol{q}(\boldsymbol{w})|\times|C_{0}|$ pick, for every $j\leq k$ with $x_{j}>0$, every $\Diamond_{i}C\in{\boldsymbol{t}}_{j}$ and every run $\rho\in\mathfrak{R}$ with $\rho(\boldsymbol{w})=j$, a successor $\boldsymbol{w}iw\in W$ with $C\in{\boldsymbol{t}}_{\rho(\boldsymbol{w}iw)}$ or $\Diamond_{i}C\in{\boldsymbol{t}}_{\rho(\boldsymbol{w}iw)}$. Let $V$ be the set of successors of $W$ picked in this way. Then construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by restricting $\mathfrak{Q}$ to
$W^{\prime}=(W\setminus W_{\boldsymbol{w}})\cup\{\boldsymbol{w}\}\cup\bigcup_{\boldsymbol{v}\in V}W_{\boldsymbol{v}}.$
It is easy to see that $\mathfrak{Q}^{\prime}$ is a quasimodel for $C_{0}$ such that the outdegree of $\boldsymbol{w}$ is bounded by $|\boldsymbol{q}(\boldsymbol{w})|\times|C_{0}|$.
@@
As argued above, it follows from Lemma 28 that we have a recursive bound on the size of a quasimodel for $C_{0}$. By Lemma 27, we then also have a recursive bound on the size of a finite tree-shaped model satisfying $C_{0}$. This shows the decidability of concept satisfiability for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$ in expanding domain models.
D.3 Noetherian Frames
We now discuss concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ and $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ in models with expanding domains. We focus on $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$; the case of $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ can be treated similarly. Note that a straightforward selective filtration argument applied to appropriate quasimodels shows that every $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$-concept satisfiable in a model with expanding domains is satisfied in a finite model with expanding domains. This observation implies that a $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}}$-concept $C$ with a single modal operator is satisfiable in an expanding domain model for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ iff it is satisfiable in a model for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ with expanding domains, where the modal operator in $C$ is interpreted by the transitive closure of $R_{1}\cup\cdots\cup R_{n}$. Hence the decidability of $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$-concept satisfiability in expanding domain models is a consequence of the decidability of $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$-concept satisfiability in expanding domain models.
The decidability of concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ in expanding domain models can also be derived from decidability results for expanding domain product modal logics. We define the relevant modal logics using DL syntax. Consider the modal DL with a single role, $o$, interpreted as a Noetherian strict linear order. In DL syntax, concepts in $\mathcal{ML}^{n}_{lin}$ are of the form
@@ D.3 Noetherian Frames
The proof is similar to the proof of Theorem 10 (1). Assume an $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concept $C$ is given. Assume for simplicity that $C$ contains a single modal operator, denoted by $\Box$. Let $m$ be the modal depth of $C$. Let $\mathsf{con}(C)$ be the closure under single negation of the set of concepts occurring in $C$. A type for $C$ is a subset ${\boldsymbol{t}}$ of $\mathsf{con}(C)$ such that $\neg C\in{\boldsymbol{t}}$ iff $C\not\in{\boldsymbol{t}}$, for all $\neg C\in\mathsf{con}(\mathcal{O})$. A quasistate for $C$ is a non-empty set $\boldsymbol{T}$ of types for $C$. The description of $\boldsymbol{T}$ is the $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concept
@@
Now decidability of $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$-concept satisfiability in expanding domain models follows from decidability of satisfiability of $\mathcal{ML}^{1}_{lin}$-concepts in models with expanding domains and transitive Noetherian frames (?).